OpenStudy (anonymous):

Use the Gauss's approach to find the following sums. A. 1+2+3+4+...+1001 B. 1+3+5+7+...+101 The sum of the sequence is? The sum of the sequence is?

OpenStudy (amistre64):

the gauss approach is about 6 feet under i think

OpenStudy (anonymous):

gauss's approach (so legend has it) is to add forwards and backwards and then divide by 2

OpenStudy (kinggeorge):

A. 1+2+3+4+...+1000+1001 = (1+1000)+(2+999)+(3+998)+...+(500+501)+1001

OpenStudy (anonymous):

\[S=1+2+3+..+1000+1001\] \[S=1001+1000+...+3+2+1\] \[2S=1001\times 1002\] \[S=\frac{1001\times 1002}{2}\]

OpenStudy (kinggeorge):

From that we have 1001+1001+1001+...+1001+1001 exactly 501 times. So take \[1001\cdot500+1001 = 500500+1001=501501\]

OpenStudy (anonymous):

Thanks for showing me how to work it. I can work the other problems like these one now.

OpenStudy (kinggeorge):

I can show you how to do B also if you want me to.

OpenStudy (anonymous):

Yes thats fine

OpenStudy (kinggeorge):

As for B\[1+3+5+...+99+101=(1+99)+(3+97)+...+(49+51)+101\]Which is\[100+100+100+...+101\]With 25 terms equal to \(100\) This leaves us with \[25\cdot100 +101=2500+101=2601\]

OpenStudy (anonymous):

Thanks KingGeorge

OpenStudy (kinggeorge):

You're welcome.